OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)= 8/x g(x)= 8/x

OpenStudy (anonymous):

\[g(f(x))=g(\frac{8}{x})=\frac{8}{\frac{8}{x}}\] is a start

OpenStudy (anonymous):

plug in x = g(x) into f(x)

OpenStudy (anonymous):

how can i do that for this one?

OpenStudy (anonymous):

g(x) = 8/x while f(x) = 8/x, plugging in x = g(x) into f(x) you have 8/(8/x) or 8 * x/8

OpenStudy (anonymous):

that's how it's confirmed?

OpenStudy (anonymous):

Simplify the fraction and look at the question again.. youll see it

OpenStudy (anonymous):

we have to show that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

yes, if you look at the first step that i wrote, you get \(g(f(x))=x\) in one more step